# XY-Cut++: Advanced Layout Ordering via Hierarchical Mask Mechanism on a Novel Benchmark

### v1

@@ 1 Introduction
Document Reading Order Recovery is a fundamental task in document image understanding, serving as a cornerstone for enhancing Retrieval-Augmented Generation (RAG) systems and enabling high-quality data preprocessing for large language models (LLMs). Accurate layout recovery is essential for applications such as digital publishing and knowledge base construction. However, this task faces several significant challenges: (1) complex layout structure (e.g., multi-column layout, nested text boxes, non-rectangular text regions, cross-page content), (2) inefficient cross-modal alignment due to the high computational costs of integrating visual and textual features, and (3) the lack of standardized evaluation protocols for block-level reading order. Traditional approaches like XY-Cut [2] fail to model semantic dependencies in complex designs, while deep learning methods such as LayoutReader [13] suffer from prohibitive latency, limiting real-world deployment. Compounding these issues, existing datasets like ReadingBank [13] focus on word-level sequence annotations, which inadequately evaluate block-level structural reasoning—a necessity for real-world layout recovery. Although OmniDocBench [8] recently introduced block-level analysis support, its coverage of diverse complex layouts (e.g., newspapers, technical reports) remains sparse, further hindering systematic progress.
@@
    We establish a new benchmark dataset for layout ordering by curating complex layouts from document detection datasets, enriching the diversity of challenging layouts and offering a more comprehensive resource than existing datasets.
@@
    We achieved state-of-the-art results on both existing and new datasets, and provided extensive ablations to each component in our methods.
Figure 2: XY-Cut Recursive Partitioning Workflow and Failure Analysis in Complex Layouts: (1) Initial document segmentation steps, (2) Connectivity assumption violations caused by cross-layout cell structures (cell 5), and (3) Error propagation through subsequent layout ordering.
@@ 2 Related Work
Document Reading Order Recovery has seen significant advancements, evolving from early heuristic rule-based approaches to sophisticated deep learning approaches.
@@ 2.1 Traditional Approaches
Figure 3: Challenges posed by L-shaped inputs: (1) segmentation failure due to the inability to process L-shaped structures, and (2) missegmentation caused by improper handling of L-shaped regions. ( The correct order is ➁+➂ ➀)
@@ 3 Methodology
Our geometry-semantic fusion pipeline, as shown in Fig. 4, consists of four coordinated stages: (1) PP-DocLayout [10] extracts visual features and shallow semantic labels from document images; (2) highly dynamic elements (e.g., titles, tables, figures) are pre-masked to alleviate the ”L-shape” problem; (3) cross-layout elements are identified through global layout analysis, followed by mask processing, real-time density estimation, and heuristic sorting to ensure logical content ordering; and (4) masked elements are re-mapped using nearest IoU edge-weighted margins. By integrating geometry-aware processing with shallow semantic information, our method achieves state-of-the-art performance on DocBench-100 and OmniDocBench [8], demonstrating its effectiveness in addressing complex document layout challenges.
Figure 4: End-to-End Layout Ordering for Diverse Document Layouts Framework Overview: (a) Layout Detection(We use PP-Doclayout[10]), (b) Pre-Mask Processing, (c) Multi-Granularity Segmentation, and (d) Cross-Model Matching.
Figure 5: Multi-Granularity Segmentation: (1) cross-layout element masking, (2) preliminary segmentation via pre-cut, and (3) ecursive density-driven partitioning. The enhanced XY-Cut algorithm adapts its splitting axis selection through real-time density evaluation ($\tau_{d}$), prioritizing horizontal splits for content-dense regions and vertical splits otherwise.
@@ 3.1 Multi-Granularity Segmentation
Phase 1: Cross-Layout Detection Compute document-level median bounding box length and establish adaptive threshold with scaling factor $\beta=1.3$:
@@
where $c_{i}$ denotes bounding box center coordinates, and classified visual elements include {Figure, Table}. Here,$\phi_{\text{text}}(B_{i})$ represents the distance from the bounding box $B_{i}$ to the nearest text box,and $\infty$ indicates that $B_{i}$ is not adjacent to any text box.
@@
$\tau_{d}=\frac{\sum_{k=1}^{K_{c}}w_{k}^{(C_{c})}h_{k}^{(C_{c})}}{\sum_{k=1}^{K_{s}}w_{k}^{(C_{s})}h_{k}^{(C_{s})}}$ (4)
@@ 3.2 Cross-Modal Matching
Figure 7: Visualization of Complex Page $D_{c}$ from DocBench-100 Dataset Using Different Layout Analysis Methods. (a) Input Image, (b) XY-Cut[2], classic projection-based segmentation, (c) MinerU[12], an end-to-end Document Content Extraction tool, (d) XY-Cut++, our proposed method.
@@
Algorithm 1 The overall flow of cross-model matching
@@
2:Initialize Label Priority: doc title ¿ title ¿ otherwise
@@ 4 Experiments
Figure 8: Visualization of Regular Page $D_{r}$ from DocBench-100 Dataset Using Different Layout Analysis Methods. (a) Input Image, (b) XY-Cut[2], a classic projection-based segmentation, (c) MinerU[12], an end-to-end Document Content Extraction tool, (d) XY-Cut++, our proposed method.
@@ 4.1 Dataset
We introduce DocBench-100, a novel benchmark dataset constructed from the document detection dataset [10], to address the lack of a standardized dataset for evaluating Document Reading Order Recovery tasks. Existing benchmarks fail to adequately cover the diverse layout complexities encountered in real-world documents. DocBench-100 consists of two subsets: the complex subset ($D_{c}$) and the regular subset ($D_{r}$), reflecting varying levels of layout complexity found in everyday documents.
Complex Subset ($D_{c}$): This subset contains 30 pages sourced from newspapers, books, and textbooks. The column distribution is 3.3% single-column, 6.7% double-column, 90% multi-column layouts (with $\geq 3$ columns and irregular document titles). These features make $D_{c}$ particularly challenging and representative of complex, multi-column layouts typical of newspapers and scientific journals.
Regular Subset ($D_{r}$): This subset contains 70 pages sourced from academic papers, books, textbooks, exam papers, slides, financial reports, notebooks, and magazines. The column distribution is 38.6% single-column, 54.4% double-column, and 7% three-column layouts. This subset represents more common and simpler layouts typically found in academic papers and books.
Dataset Composition: To reflect the prevalence of simpler layouts in daily use, we set the ratio of $D_{r}$ to $D_{c}$ at 7:3. The entire dataset comprises 100 original page images, each accompanied by a ground truth (GT) JSON file and an input JSON (No Index) file. The GT files contain essential information, including:
  • page_id: A unique identifier for each page.
  • page_size: The dimensions of the page.
  • bbox: Bounding boxes for each element on the page.
  • label: Class labels for different types of content (e.g., text, figure).
  • index: The reading order index for each element.
All reading orders have been manually verified to ensure their accuracy.
Necessity of DocBench-100: The introduction of DocBench-100 fills a critical gap in the field by providing a comprehensive dataset that covers both simple and complex layouts. This diversity allows for robust evaluation of reading order recovery models across various scenarios. The inclusion of detailed ground truth (GT) annotations facilitates direct testing and comparison with other models, thereby promoting reproducibility and advancing further research in this domain.
@@ 4.3.1 Progressive Component Ablation
  • +Mask: Improves BLEU-4 by 6.9 points (from 0.749 to 0.818) on $D_{c}$ via adaptive thresholding (Equation (1), $\beta=1.3$), reducing false splits by 15.9%.
@@ 4.3.2 Architectural Analysis
Pre-Mask Processing (Pre-Mask): To alleviate the ”L-shaped” problem, we applied preliminary masking on highly dynamic elements such as titles, tables, and figures. This approach reduced visual noise and improved reading order recovery, resulting in a 2.5-point BLEU-4 score increase, as shown in Table 2.
Multi-Granularity Segmentation (MGS): Mask cross-layout is a very direct method to solve the problem that XY-Cut cannot segment L-shaped input. Pre-Cut realizes preliminary sub-page division through page analysis, thereby avoiding page content mixing affecting sorting. The adaptive splitting strategy enables reasonable segmentation through real-time density estimation. Table 2 shows additive benefits of the mask cross-layout (+8.3), Pre-Cut (+9.2), and the adaptive splitting scheme (+10.1) over the baseline (82.2 BLEU-4).
@@ 4.3.3 Benchmark Comparison
As shown in Table 3, our approach establishes new benchmarks across all evaluation dimensions. Specifically, it outperforms XY-Cut by a significant margin, achieving a +23.7 absolute improvement on $D_{c}$ (from 74.9 to 98.6). Additionally, it surpasses LayoutReader by +5.3 on $D_{r}$ (from 94.6 to 98.9), despite not using any learning-based components. Furthermore, our method achieves a Kendall’s $\tau$ of 0.996 overall, indicating near-perfect ordinal consistency (with $p<0.001$ in the Wilcoxon signed-rank test). Visual results presented in Fig. 7 and 8 further demonstrate the robustness in handling multi-column layouts and cross-page elements, where previous methods frequently fail.
@@ 5 Conclusion
In this work, we address the critical challenges in document layout recovery and semantic conversion, particularly focusing on the needs of Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs) for accurate information extraction. We introduce a hierarchical document parsing framework that incorporates a Mask mechanism-based layout recovery algorithm. This framework enables precise reading order recovery through pre-mask processing, multi-granularity segmentation, and cross-modal matching. Additionally, it achieves superior performance on the DocBench-100 and OmniDocbench datasets, outperforming baseline methods and providing robust support for downstream RAG and LLM tasks. It is noteworthy that while traditional methods are known for their computational efficiency, they often fall short in performance. Our approach not only maintains simplicity and efficiency but also matches or surpasses the performance of deep learning methods. This advantage is primarily attributed to two key innovations: (1) the strategic use of shallow semantic labels as structural priors during layout analysis, and (2) a hierarchical mask mechanism that effectively captures the document’s topological structure. These innovations not only enhance performance but also maintain model efficiency, suggesting promising directions for improving deep learning models in layout recovery tasks. We believe our work can inspire new approaches to tackle the challenging tasks in document understanding. Our code will be released at:https://github.com/liushuai35/PaddleXrc.
@@ 6 Limitations and Future Work
Figure 1: XY-Cut++ Architecture with Hierarchical Mask Mechanism and DocBench-100 Benchmark. (left) DocBench-100 dataset composition (9 document categories, Multi-type layout) and dual evaluation protocols: end-to-end image-based assessment (requiring layout parsing models) and direct JSON-based metric computation. (right) Algorithm workflow integrating adaptive pre-mask processing, multi-granularity segmentation, and cross-modal matching. Unlike neural network models, our method is simple, effective, and strongly interpretable.

### v2

@@ Abstract
keywords:
Document reading order, Document layout analysis, XY-Cut++, DocBench-100
Figure 1: XY-Cut++ Architecture with Hierarchical Mask Mechanism and DocBench-100 Benchmark. (left) DocBench-100 dataset composition and dual evaluation protocols. (right) Algorithm workflow integrating adaptive pre-mask processing, multi-granularity segmentation, and cross-modal matching.
@@ 1 Introduction
Document Reading Order Recovery is a fundamental task in document image understanding, serving as a cornerstone for enhancing Retrieval-Augmented Generation (RAG) systems and enabling high-quality data preprocessing for large language models (LLMs). Accurate layout recovery is essential for applications such as digital publishing and knowledge base construction. However, this task faces several significant challenges: (1) complex layout structure (e.g., multi-column layout, nested text boxes, non-rectangular text regions, cross-page content), (2) inefficient cross-modal alignment due to the high computational costs of integrating visual and textual features, and (3) the lack of standardized evaluation protocols for block-level reading order. Traditional approaches like XY-Cut [1] fail to model semantic dependencies in complex designs, while deep learning methods such as LayoutReader [2] suffer from prohibitive latency, limiting real-world deployment. Compounding these issues, existing datasets like ReadingBank [2] focus on word-level sequence annotations, which inadequately evaluate block-level structural reasoning—a necessity for real-world layout recovery. Although OmniDocBench [3] recently introduced block-level analysis support, its coverage of diverse complex layouts (e.g., newspapers, technical reports) remains sparse, further hindering systematic progress. Meanwhile, large-scale layout datasets such as DocBank [4] and PubLayNet [5] have advanced document layout analysis but primarily target word/line-level or scientific article layouts and do not directly benchmark block-level reading order.
@@
    We curate DocBench-100, a block-level benchmark with diverse complex layouts that complements existing datasets and directly targets reading-order evaluation.
@@
    We achieved state-of-the-art results on both existing and new block-level datasets, and provided extensive ablations to each component in our methods.
Figure 2: XY-Cut Recursive Partitioning Workflow and Failure Analysis in Complex Layouts: (1) Initial document segmentation steps, (2) Connectivity assumption violations caused by cross-layout cell structures (cell 5), and (3) Error propagation through subsequent layout ordering. The correct reading order is 1, 3, 2, 4, 5, 6, 7.
@@ 2 Related Work
Document Reading Order Recovery has advanced from early heuristic rule-based methods to modern deep learning models, yet reliably handling complex layouts remains challenging.
@@ 2.1 Traditional Approaches
Figure 3: Challenges posed by L-shaped inputs: (1) segmentation failure due to the inability to process L-shaped structures, and (2) missegmentation caused by improper handling of L-shaped regions. The correct sequence of segmentation is ➁+➂ ➀.
@@ 2.3 Benchmarks for Document Reading Order Recovery
3 DocBench-100 Dataset
To enable rigorous evaluation of block-level reading order, we construct DocBench-100, a curated benchmark emphasizing diverse real-world layouts. It complements existing datasets (ReadingBank [2], OmniDocBench [3], DocBank [4], PubLayNet [5]) by focusing on page-level block ordering across complex structures.
3.1 Sources and Composition
We source candidate pages from public document detection corpora (notably PP-DocLayout [19]) and MinerU [20] extraction outputs, then select pages exhibiting challenging phenomena: multi-column articles, irregular or spanning titles, interleaved figures/captions, and nested regions. The dataset comprises 100 pages split into: complex ($D_{c}$, 30 pages) and regular ($D_{r}$, 70 pages), following the prevalence in practical applications. An overview will be presented in Fig. 4.
3.2 Fields and File Structure
Each page includes an image and two JSON files: (1) an input JSON (No Index) with fields page_id, page_size, bbox, label; (2) a GT JSON that additionally includes block-level reading order index. This design supports both end-to-end and oracle-detection evaluation protocols.
3.3 Annotation Pipeline
We adopt a two-stage pipeline aligned with reviewer suggestions:
  • Automatic pre-annotation: MinerU [20] provides initial blocks, labels, and an order hypothesis.
  • Manual verification and screening: Annotators correct segmentation and labels, and assign final index. Pages with inherently ambiguous global reading order (e.g., collage-like scans lacking semantic anchors) are excluded. All remaining reading orders are manually verified.
3.4 Statistics and Usage Protocols
$D_{c}$ contains predominantly multi-column layouts ($\geq 3$ columns) and irregular titles, while $D_{r}$ is dominated by single/double columns common in academic and business documents. Summary statistics by column structure are reported in Table 1. We recommend reporting both: (a) end-to-end image-based evaluation (requires detection) and (b) JSON-based evaluation on block sequences. We will release the evaluation script for block-level BLEU and dataset documentation.
Table 1: DocBench-100 subset statistics by column structure. Percentages are computed within each subset.
Subset Pages Single Double $\geq 3$
$D_{c}$ 30 3.3% 6.7% 90.0%
$D_{r}$ 70 38.6% 54.4% 7.0%
Figure 4: DocBench-100 image overview: (a) 70-page regular subset $D_{r}$, and (b) 30-page complex subset $D_{c}$. All pages provide block-level reading-order ground truth for benchmarking layout-ordering methods.
@@ 4 Methodology
Our geometry–semantic fusion pipeline (Fig. 5) has four stages: (1) PP-DocLayout [19] extracts visual features and shallow semantic labels; (2) highly dynamic elements (e.g., titles, tables, figures) are pre-masked to alleviate the L-shape problem; (3) cross-layout elements are identified via global analysis, then processed with masking, real-time density estimation, and heuristic sorting; and (4) masked elements are re-mapped using nearest IoU edge-weighted margins. This geometry-aware pipeline with shallow semantics attains state-of-the-art results on DocBench-100 and OmniDocBench [3].
Figure 5: End-to-End Layout Ordering for Diverse Document Layouts Framework Overview: (a) Layout Detection ($P$-DocLayout [19]), (b) Pre-Mask Processing, (c) Multi-Granularity Segmentation, and (d) Cross-Modal Matching.
Figure 6: Multi-Granularity Segmentation: (1) cross-layout element masking, (2) preliminary segmentation via pre-cut (①), and (3) recursive density-driven partitioning. The enhanced XY-Cut algorithm adapts its splitting axis selection through real-time density evaluation ($\tau_{d}$), prioritizing horizontal splits for content-dense regions and vertical splits otherwise (②③).
@@
4.1 Pre-Mask Processing
We first identify candidate high-dynamic elements using shallow semantic labels provided by PP-DocLayout [19] and simple geometric heuristics (e.g., large area, extreme aspect ratio, or spanning multiple columns). We then construct a binary mask that temporarily excludes these elements from the subsequent multi-granularity segmentation and geometric sorting stages. During training, we randomly perturb the mask to increase robustness to imperfect detection and to teach the sorting heuristics to infer global structure without relying on a few dominant blocks. At inference time, masked elements are re-mapped back to the ordered layout using a lightweight cross-modal matching strategy: candidate masked blocks are matched to nearby unmasked regions by an IoU and margin-weighted nearest-neighbor rule that favors semantic priority (title $\succ$ text $\succ$ figure). This two-stage handle—mask first then remap—reduces L-shape artifacts while preserving final layout fidelity.
@@ 4.2 Multi-Granularity Segmentation
Phase 1: Cross-Layout Detection Compute document-level median bounding box length and establish an adaptive threshold with scaling factor $\beta=1.3$:
@@
[…]
@@
$\tau_{d}=\frac{\sum_{k=1}^{N_{c}}w_{k}^{(C_{c})}h_{k}^{(C_{c})}}{\sum_{k=1}^{N_{s}}w_{k}^{(C_{s})}h_{k}^{(C_{s})}}$ (4)
where $C_{c}$ denotes cross-layout elements and $C_{s}$ represents single-layout elements within region $R$. $N_{c}$ and $N_{s}$ are their respective counts.
@@ 4.3 Cross-Modal Matching
Figure 8: Visualization of Complex Page $D_{c}$ from DocBench-100 Dataset Using Different Layout Analysis Methods.
(a) Input Image, (b) XY-Cut[1], classic projection-based segmentation, (c) MinerU[20], an end-to-end Document Content Extraction tool, (d) XY-Cut++, our proposed method.
Figure 9: Visualization of Regular Page $D_{r}$ from DocBench-100 Dataset Using Different Layout Analysis Methods.
(a) Input Image, (b) XY-Cut[1], a classic projection-based segmentation, (c) MinerU[20], an end-to-end Document Content Extraction tool, (d) XY-Cut++, our proposed method.
@@
Algorithm 1 The overall flow of cross-modal matching
@@
2:Initialize Label Priority: $\text{doc title}\succ\text{title}\succ\text{otherwise}$
@@ 5.1 Dataset
We evaluate on DocBench-100; detailed sources, fields, annotation pipeline, screening criteria, statistics, and usage protocols are described in Section 3. For fairness, we report both end-to-end and JSON-based protocols, and clarify FPS timing scope alongside Table 6.
@@ 5.2 Setup
Block-level BLEU (detection order). We evaluate the ordering of detection boxes (block IDs), not textual content. Following the original BLEU definition [21], we compute n-gram precisions on block identifiers and report BLEU-4 with brevity penalty:
$\mathrm{BLEU}\_4(\hat{s},s)=BP\cdot\exp\left(\sum_{n=1}^{4}\frac{1}{4}\log p_{n}\right),$ (11)
where $p_{n}$ is the precision of block-level n-grams and $BP=\begin{cases}\exp(1-r/c),&c\leq r\\ 1,&c>r\end{cases}$ with $c$ the hypothesis length and $r$ the reference length.
FPS measurement scope. Table 6 reports FPS for the ordering/sorting module only; upstream detection (e.g., PP-DocLayout) and downstream OCR/LM are excluded. Results are averaged over 10 runs on Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz, 256GB RAM.
@@ 5.3.1 Progressive Component Ablation
  • +Pre-Mask: Improves BLEU-4 by 6.9 points (from 0.749 to 0.818) on $D_{c}$ via adaptive thresholding (Equation (1), $\beta=1.3$), reducing false splits by 15.9%.
@@ 5.3.2 Architectural Analysis
Pre-Mask Processing (Pre-Mask): To alleviate the ”L-shaped” problem, we applied preliminary masking on highly dynamic elements such as titles, tables, and figures (see Methodology Section 4.1). This approach reduced visual noise and improved reading order recovery, resulting in a 2.5-point BLEU-4 score increase, as shown in Table 3.
extbfMulti-Granularity Segmentation (MGS): Mask cross-layout is a very direct method to solve the problem that XY-Cut cannot segment L-shaped input. Pre-Cut realizes preliminary sub-page division through page analysis, thereby avoiding page content mixing affecting sorting. The adaptive splitting strategy enables reasonable segmentation through real-time density estimation. Table 3 shows additive benefits of the mask cross-layout (+8.3), Pre-Cut (+9.2), and the adaptive splitting scheme (+10.1) over the baseline (0.822 BLEU-4).
@@ 5.3.3 Benchmark Comparison
As shown in Table 4, our approach establishes new benchmarks across all evaluation dimensions. Specifically, it outperforms XY-Cut by a significant margin, achieving a +23.7 absolute improvement on $D_{c}$ (from 74.9 to 98.6). Additionally, it surpasses LayoutReader by +5.3 on $D_{r}$ (from 94.6 to 98.9), despite not using any learning-based components. Furthermore, our method achieves a Kendall’s $\tau$ of 0.996 overall, indicating near-perfect ordinal consistency (with $p<0.001$ in the Wilcoxon signed-rank test). Visual results presented in Fig. 9 and 9 further demonstrate the robustness in handling multi-column layouts and cross-page elements, where previous methods frequently fail.
@@ 6 Conclusion
In this work, we address the critical challenges in document layout recovery and semantic conversion, particularly focusing on the needs of Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs) for accurate information extraction. We introduce a hierarchical document parsing framework that incorporates a Mask mechanism-based layout recovery algorithm. This framework enables precise reading order recovery through pre-mask processing, multi-granularity segmentation, and cross-modal matching. Additionally, it achieves superior performance on the DocBench-100 and OmniDocBench datasets, outperforming baseline methods and providing robust support for downstream RAG and LLM tasks. It is noteworthy that while traditional methods are known for their computational efficiency, they often fall short in performance. Our approach maintains simplicity and efficiency while matching or surpassing deep learning methods. This advantage stems from two innovations: (1) shallow semantic labels used as structural priors, and (2) a hierarchical mask mechanism that captures document topology. These improve performance without sacrificing efficiency and suggest promising directions for future work. Code will be released at: https://github.com/liushuai35/PaddleXrc.git.
@@ 7 Limitations and Future Work
8 Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Additional Visual Results
A.1 Other Cases
Figure 10: DocBench-100 (complex subset) — Example A. A challenging multi-column page with spanning titles and interleaved elements. XY-Cut++ maintains a coherent block-level reading order across columns by combining pre-mask and density-aware splitting.
Figure 11: DocBench-100 (complex subset) — Example B. Nested regions and interleaved figures/captions. The mask-then-remap strategy reduces L-shape artifacts and preserves local grouping for globally consistent ordering.
Figure 12: DocBench-100 (complex subset) — Example C. L-shaped text around graphics. Pre-mask normalization avoids early incorrect splits, and the final sequence aligns with GT ordering.
Figure 13: DocBench-100 (complex subset) — Example D. Newspaper-like dense columns. The adaptive axis selection (horizontal vs. vertical) driven by regional density yields stable column-wise ordering.
Figure 14: DocBench-100 (regular subset) — Example A. A standard single/double-column page illustrating the method’s stability on common business and academic layouts.
Figure 15: OmniDocBench (complex) — Example A. A multi-column page with captions and side notes; geometry-aware matching preserves inter-column flow and caption linkage, illustrating robustness on challenging document layouts.
Figure 16: OmniDocBench (double) — Example B and ancient-text case. Despite right-to-left script conventions and sparse annotations, the method maintains correct block-level order, evidencing robustness to atypical typography.
A.2 Failure Cases
Figure 17: Challenging cases on OmniDocBench and DocBench-100. Failures stem from (1) insufficient fine-grained semantics and (2) sub-page complexity, leading to sorting errors; these motivate sub-page detection and stronger semantic priors.